OpenStudy (anonymous):

At what values of x is f(x)=(x^2-2x-3)/(x-2)=0 ?

OpenStudy (angela210793):

f(x)=(x^2-2x-3)/(x-2)=0 when x^2-2x-3=0 so D=b^2-4ac=4-4*1*(-3)=16 x1=(-b+sqrD)/2a==(2+4)2=3 x2=(-b-sqrD)/2a=(2-4)/2=-2

OpenStudy (anonymous):

\[f(x) = \frac{x^2-2x-3}{x-2}=0\] When \[x^2-2x-3=0\] (Denominator can't make that equation 0; only numerator can.) Factor that polynomial, and you will get:\[(x-3)(x+1)=0\] And its zeros are therefore: \[x = 3, -1\] If you try to plug in 3, you will get:\[\frac{3^2-2(3)-3}{3-2}=0\] Do the same for -1: \[\frac{(-1)^2-2(-1)-3}{3-(-1)}=0\] Do be mindful that if a solution makes the denominator 0, then that solution yields an undefined value.

OpenStudy (angela210793):

yea -1 sorry not -2 (typing mistake) :(